Colored Wheel



A survey on 200 Senior High School learners on how much time they read per



week has a confidence interval of (13.5, 17.9) in hours at 95% confidence level.



What is the margin of error? _________________________



What is the average time they read per week? __________________


1
Expert's answer
2022-03-15T03:28:15-0400

Since the confidence interval is symmetric,

The margin of error is "\frac{17.9-13.5}{2}=2.2"

The average time is "\frac{13.5+17.9}{2}=15.7"